# Cross-Calibration and First Vertical ECE Measurement of Electron Energy Distribution in the TCV Tokamak

### v1

@@ 1 Introduction
Electron Cyclotron Emission (ECE) is a well-established technique for measuring electron temperature in Tokamaks [1, 2]. The conditions for such measurement generally include a thermal plasma, i.e., a plasma in which the electron distribution can be assumed Maxwellian, and optically thick conditions for the electron cyclotron wave. For temperature measurement, when these conditions are fulfilled, ECE configurations are usually chosen to have a horizontal line of sight, viewing the plasma perpendicularly from the Low Field Side (LFS ECE) or from the High Field Side (HFS ECE). In these horizontal configurations, ECE measurements at the $n$ harmonic take advantage of the $1/R$ dependence of the electron cyclotron frequency, $\omega_{n\mathrm{ece}}\sim neB(R)/m_{\mathrm{e}}$, to infer the radial profile of the electron temperature, $T_{\mathrm{e}}(R)$. Electron Cyclotron Emission has also been used to diagnose plasmas in which the electron distribution has departed from a Maxwellian distribution. For the diagnosis of non-thermal electrons, the Vertical ECE configuration (V-ECE), with $R$ constant along the line of sight, has proven to have the advantage of a more straightforward discrimination of the electron energy according to the radiation frequency:
@@
The work presented in this paper, on the analysis of the electron energy distribution using V-ECE measurements, builds on previous works on Alcator C [3, 4, 5, 6, 7, 8], TEXT-U [9, 10], WT-3 [11], DIII-D [12, 13], and PLT [14, 15, 16] Tokamaks, which attempted to leverage the ideal one-to-one correspondence between measured frequency and electron energy in Equation 1, by minimizing or mitigating issues such as harmonic overlap and multiple wall reflections. A method, described in Reference [17], was also attempted to infer electron temperature from V-ECE measurements. In the following sections, we will discuss the experimental arrangement that enables the measurement of non-thermal electrons using V-ECE on TCV (Section 2). The calibration of the diagnostic, which relies on V-ECE modeling and measurements free of background radiation, will be presented in Section 3. V-ECE measurements of fast electron radiation in runaway electron scenarios and Electron Cyclotron Current Drive (ECCD) will be discussed in Section 4. The analysis of the electron energy distribution based on these measurements is presented in Section 5, followed by the general conclusion of the paper.
@@ 2 Experimental Setup
The measurements discussed in this paper are carried out using the V-ECE diagnostic[18], installed on the TCV Tokamak [19]. The diagnostic is equipped with a set of heterodyne radiometers measuring in the frequency range $78-114$ GHz, with temporal resolution in the order of $10\mu$s and bandwidth $\sim 750$ MHz. The measured radiation is split into the X- and O- polarisation (X- and O-mode) using a wire grid polariser. The optics of the diagnostic is arranged such that the vertical antenna pattern forms a Gaussian beam of waist radius $\sim 3$ cm near the vessel mid-plane. In vacuum, the antenna pattern terminates, at the bottom of the machine, on a highly absorbing viewing dump [20], see Figure 1a. During plasma discharges, the antenna pattern can be shifted away from the viewing dump, due to refraction, at a threshold electron density close to $1\times 10^{19}\mathrm{m}^{-3}$[21]. The diagnostic measures fast electron emission at downshifted third or fourth harmonic emission as shown in Figure 1b. The electron energy that can be measured at each frequency is calculated in[18]. It is calculated that the maximum energy the diagnostic can measure without harmonic overlap is in the order of $\sim 250$ keV. The ECCD experiments discussed in this paper exploit the X2 gyrotron from the ECH/ECCD system on TCV[22]. The gyrotron produces wave at $82.7$ GHz and nominal power $\sim 680$ kW. The gyrotron is connected to a launcher which direct the beam into the plasma. Current drive is achieved by imparting a parallel component, with respect to the plasma magnetic field, to the ECH wave vector. The location of the beam power deposition along the radial coordinate is generally controlled by the chosen value of magnetic field strength. The toroidal and poloidal directions of the beam ((((in the reference frame of the plasma )))) are controlled by the launcher angles, so-called launcher toroidal angle, $\phi_{\mathrm{L}}$, and poloidal angle, $\theta_{\mathrm{L}}$ that can can be varied during discharges.
Figure 1: 1a V-ECE optics on TCV. The line of sight, which terminates on a beam dump (dark component below the machine floor) forms a Gaussian beam of radius $\sim$ 3 cm near the mid-plane. The radiation is reflected on an ellipsoidal mirror, split into X- and O- polarisations and detected in a heterodyne radiometer. 1b V-ECE frequency range and ECE harmonics on TCV with a toroidal magnetic field $B_{0}$=1.42 T at the vacuum vessel centre $R_{0}$=0.88 m The represented plasma frequency is computed with a central electron density of 1.5 ×$\times 10^{19}$ m -3.
@@ 3.1 Background
The calibration of ECE diagnostics generally consists in determining, with the highest accuracy, the spectral response of the system to a given radiation power. Traditionally, two approaches have been retained for ECE calibration: (1) the use of sources of known characteristics as substitutes for the plasma and (2) the cross-calibration on a different, absolutely calibrated instrument. In the former approach, hot and cold sources are positioned within the vacuum chamber to irradiate the ECE antenna. The approach has been used since the seventies in the vast majority of ECE systems in magnetic confinement devices [23, 24]. A limitation of this approach is that the calibration needs to be conducted inside the vacuum vessel, necessarily during a machine vent. The system’s response, determined during a hot/cold source calibration, is then used for several months, or even years, until a subsequent calibration is attainable. This raises the question of the calibration’s stability in relation to variations in experimental conditions, which may occur on a daily basis. The latter calibration approach exploits the absolute calibration of a different instrument within the same Tokamak, to obtain the response function of an ECE system. The ECE system can be cross-calibrated on another ECE system as in JET [25], or can be cross-calibrated on a different diagnostic which measures the same plasma parameters such as Thomson Scattering, inheriting all the uncertainties from the absolute calibration. This calibration method relies on the assumption that both diagnostics measure the same electron temperature. For ECE diagnostic systems, the measurement of the electron temperature requires that the plasma is optically thick and that the electron distribution is Maxwellian. When these conditions are fulfilled, successful cross-calibration of ECE systems with Thomson Scattering (TS) can be achieved, as demonstrated over the years, on TCV for example [26, 27, 28, 29, 30]. Historically, ECE and Thomson Scattering cross-calibration have used ohmic plasma discharges, which provide the best experimental plasma conditions allowing the assumption of a Maxwellian electron distribution. In other experimental conditions, such as those with NBI or ICRH heating, discrepancies between ECE and Thomson Scattering can occur, and modeling of the electron distribution is required [31, 32]. For a vertical ECE configuration, a direct cross-calibration with Thomson Scattering cannot be achieved for two main reasons. First, even in an optically thick plasma, the radiation intensity measured with a vertical line of sight cannot be straightforwardly associated with a local electron temperature, since the ECE resonance is not well localized in the vertical direction. Second, the plasma is not usually optically thick for a vertical ECE line of sight above a certain frequency (approximately 86 GHz on TCV). At these higher, optically thin frequencies, an accurate estimation of the absorption coefficient and, consequently, the radiation intensity needs to be performed in order to use the plasma emission for calibration. Examples of calculations of radiation intensity in optically thin conditions have been reported in previous works, such as in DIII-D for the measurement of the electron temperature from the optically gray third harmonic [33], and in JET for calibrating the optically thin O1 (first harmonic O-mode) radiation for temperature measurement in the inboard (High Field Side) pedestal [34]. In both cases, a horizontal lines of sight was employed.
@@ 3.4.1 Calculation of the peak power in a frequency band
The calibration factors are obtained by matching the modeled peak power with the measured voltage. The spectral intensities calculated for a few frequencies during discharge #$\#67946$ are shown in Figure 7 with their profile in each bandwidth. The parabolic shape of the intensity profile in the bandwidth justifies the need for optimization of the frequency step. The bandwidth power $P_{\mathrm{BW}}$ is calculated as:
@@ 3.6 Discussing the uncertainties
The uncertainty on the V-ECE signal comes essentially from the digital filtering of the raw data. For the case of interest, the relative uncertainty on the filtered value used for the calibration is $\delta_{V}/{V}<5\%$. To assess the uncertainty in the modeled bandwidth power, the uncertainties in the parameters are propagated from the raw measurement to the calculated bandwidth power, as shown in in Figure 9. The assessment does not account for the potential uncertainty in the measurement of the magnetic field because the diagnostic is calibrated with thermal peak measurement, when the resonance is within the line of sight. The measured frequencies and the the line of sight constrain the value of the magnetic field at the resonance. The uncertainties that are propagated are those from the measurement of the electron density, $\delta_{n_{\mathrm{e}}}$ and electron temperature, $\delta_{T_{\mathrm{e}}}$. On TCV, The relative errors for both density and temperature measurements with Thomson Scattering are respectively below $5\%$ and $10\%$ in the plasma center. Both values increase significantly at the plasma edge, up to $25\%$ and $40\%$ respectively. The high uncertainties on the density and temperature profiles at the edge do not impact significantly the modeled intensity. That is because the optical depth of the plasma drops at the edge in such a way that the intensity depends mainly on the parameters in the plasma center.
@@ 4.1 ECCD measurement
Figure 11: Raw measurements from the discharges #$\#73217$ and #$\#72644$ in the frequency range $96-114$ GHz.
@@
Calibrated V-ECE measurements during ECCD discharges #$\#73217$ and #$\#72644$ on TCV have been presented in Reference [18]. Raw measurements during those discharges are shown in Figure 11. Both discharges have similar values of ECH power. While the launcher toroidal angle is kept constant, the poloidal angle is varied in $5$ steps during the heating phase between $0.7$ s and $1.9$ s in #$\#72644$. In this configuration, as the EC beam propagates along the plasma equatorial mid-plane, the launcher angle controls the parallel component of the wave vector with respect to the toroidal magnetic field in the plasma. When this parallel wave-vector component is zero ($\phi_{\mathrm{L}}=0$), the EC wave purely heats the plasma by increasing the perpendicular energy of the resonant electrons around a given parallel velocity. Increasing the parallel wave vector (by increasing $\phi_{\mathrm{L}}$ ) results in a Doppler shift of the EC resonance, pulling a suprathermal electron tail out of the bulk at higher velocities than the corresponding heating case [40]. The variation of the poloidal angle of the launcher from $\sim 10^{\circ}$ to $\sim 26^{\circ}$ in each step allows the observation of a stair-shaped X-mode radiation intensity. When no heating power is applied in both plasma discharges, the measured intensities coincide and, are in general close to the noise level of the radiometer since no background radiation pollutes the measurement. As soon as the heating is turned on in the discharge #$\#73217$ at $\sim 0.3$ s, all the channels immediately peak to reach a nearly stationary intensity level. We note how the intensity level during the discharge #$\#73217$ is generally higher than the intensity level in the discharge #$\#72644$ when the launcher configurations are similar. This is due to the more efficient current drive achieved in #$\#73217$, due in part to the lower density and higher electron temperature as the current drive efficiency is proportional, in general, to the ratio $T_{e}/n_{e}$ [41]. When the heating is turned on during the discharge #$\#72644$, at $\sim 0.7$ s, only the low energies (((($\sim$ $5-26$ keV)))) exhibit a sharp jump in the measured intensity. The intensity at the higher energies ($\sim$ $67-75$keV) shows a similar jump, with a delay of more than $\sim 500$ ms, when the launcher angle has been moved to its third stage. From that moment, the dynamics is inverted and the higher energies now have the sharpest jump in intensity compared to the lower energies. Interestingly, the measurement at the middle energies, in our range, will then reach the highest value of intensity when the launcher angle reaches its last stage. In measurements with Hard X-Ray tomographic Spectroscopy system (HXRS) [42] during the discharge #$\#72644$, the count rates shown in Figure 12 allow similar observations on the effects of the launcher angle sweep. In TCV, hard x-rays are emitted by suprathermal electron bremsstrahlung and are measured between 20 and 200 keV with an 8 keV resolution. In Figure 12, the sampling interval is set to 10 keV, which aligns with the resolution, making finer sampling unnecessary
@@
Simulations of the ECH wave interaction with the plasma, for the discharge #$\#72644$ is achieved using the the ray tracing code C3PO, solving the linear wave power absorption [43]. The refractive indices and the angle between the wave vector and the magnetic field, calculated at the wave absorption location for the various launcher angles are shown in Table 1. The increase in the launcher poloidal angle allows to achieve higher values ((((in absolute terms)))) of the parallel refractive index. Higher parallel refractive indices generally allow the intersection between the EC resonant curves ($v_{\perp}$ vs $v_{\parallel}$) and the $v_{\parallel}$ axis at higher values of $v_{\parallel}$ [44], see Figure 13. The physical mechanisms of the plasma-wave interaction mechanisms include, in general, momentum diffusion and pitch-angle scattering, which make the ECCD resonant interaction region complex to determine. For the purposes of this experiment it is observed that higher values $\theta_{L}$ and thus $N_{\parallel}$ induce more energetic electrons in the plasma, explaining the different jumps in the measured intensities.
@@
Figure 14 show the jumps in intensity, relative to the thermal phase, measured at each frequency in the different stages of discharge #$\#72644$. The jump is obtained by dividing the intensity in each heated stage by the intensity in the thermal phase. It is seen how at relatively low parallel refractive index (((($N_{\parallel}=-0.22$)))), the highest frequencies ((((lowest energies)))) jump is the most important. As $N_{\parallel}$ increases, an intermediate frequency (((( intermediate energy as well)))), near $104$ GHz shows the sharpest jumps. We note that V-ECE X-mode intensity in ideal conditions depends on the density of the non-thermal electrons in a given energy band. That is, the highest jump observed around $104$ GHz, with increasing $N_{\parallel}$ suggests an increase in the number density of non-thermal electrons at the corresponding energy. This does not necessarily mean that the ECH wave is damped on electrons whose V-ECE energies are measured at $104$ GHz when the refractive index increases. The increase of the number density at the given energy, which explains the sharp jump, is a consequence of the overall dynamics in the electron energy space, originally caused by the ECH wave.
@@
First-principle calculation using the code LUKE is used compute the electron energy distribution at the different stages of discharge #$\#72644$. The first calculation is shown in Figure 15a. The total number of electrons above a certain energy is plotted for a Maxwellian distribution and for the computed distribution. At a given energy, say 15 keV, which is above the electron thermal energy, the total number of electrons above that energy increases over time intervals ((((with increasing $N_{\parallel}$)))). This results confirms that low-energy electrons are pushed into the fast-electrons tail with ECCD and the number of transported electrons increases with parallel refractive index. Figure 15b shows the derivative of the number of electrons in the modeled distribution deduced from the density in the Maxwellian distribution. The discontinuity in the spectra is due to the logarithmic scale and to the absolute value in the plotted quantity. It represents the regions where the density in the Maxwellian distribution changes from higher to lower than that of the modeled distribution, which is the region in the phase space where the electrons are being pumped out by the ECH wave.

### v2

@@ 1 Introduction
Electron Cyclotron Emission (ECE) is a well-established technique for measuring electron temperature in Tokamaks [1, 2]. The conditions for such measurement generally include a thermal plasma, i.e., a plasma in which the electron distribution can be assumed Maxwellian, and optically thick conditions for the electron cyclotron wave. For temperature measurement, when these conditions are fulfilled, ECE configurations are usually chosen to have a horizontal line of sight, viewing the plasma perpendicularly from the Low Field Side (LFS ECE) or from the High Field Side (HFS ECE). In these horizontal configurations, ECE measurements at the $n$ harmonic take advantage of the $1/R$ dependence of the electron cyclotron frequency [3], $\omega_{n\mathrm{ece}}\sim neB(R)/m_{\mathrm{e}}$, to infer the radial profile of the electron temperature, $T_{\mathrm{e}}(R)$. Electron Cyclotron Emission has also been used to diagnose plasmas in which the electron distribution has departed from a Maxwellian distribution. For the diagnosis of non-thermal electrons, the Vertical ECE configuration (V-ECE), with $R$ constant along the line of sight, has proven to have the advantage of a more straightforward discrimination of the electron energy according to the radiation frequency:
@@
The work presented in this paper, on the analysis of the electron energy distribution using V-ECE measurements, builds on previous works on Alcator C [4, 5, 6, 7, 8, 9], TEXT-U [10, 11], WT-3 [12], DIII-D [13, 14], and PLT [15, 16, 17] Tokamaks, which attempted to leverage the ideal one-to-one correspondence between measured frequency and electron energy in Equation 1, by minimizing or mitigating issues such as harmonic overlap and multiple wall reflections. A method, described in Reference [18], was also attempted to infer electron temperature from V-ECE measurements. In the following sections, we will discuss the experimental arrangement that enables the measurement of non-thermal electrons using V-ECE on TCV (Section 2). The calibration of the diagnostic, which relies on V-ECE modeling and measurements free of background radiation, will be presented in Section 3. V-ECE measurements of fast electron radiation in runaway electron scenarios and Electron Cyclotron Current Drive (ECCD) will be discussed in Section 4. The analysis of the electron energy distribution based on these measurements is presented in Section 5, followed by the general conclusion of the paper. All physical quantities in this paper are expressed using the International System of Units (SI). The angular frequency $\omega$ in rad/s will be preferred for equations, while the frequency $f$ in Hz will be used in quantitative assessments for practical reasons.
@@ 2 Experimental Setup
The measurements discussed in this paper are carried out using the V-ECE diagnostic[19], installed on the TCV Tokamak [20]. The diagnostic is equipped with a set of heterodyne radiometers measuring in the frequency range $78-114$ GHz, with temporal resolution in the order of $10\mu$s and bandwidth $\sim 750$ MHz. The radiometer, described in detail in [21], is composed of three main stages: a Radio Frequency (RF) stage, an Intermediate Frequency (IF) stage, and a video stage. The incoming radiation from the plasma is split into the X- and O- polarisation (X- and O-mode) using a wire grid polariser, then downshifted to lower frequencies in the RF stage. This frequency conversion occurs in a mixer, which is a diode with a non-linear I/V characteristic. The lower frequency signals are amplified in the IF stage, rectified by a square-law diode detector, and further processed in the video stage, where the signals are amplified and low-pass filtered down to 100 kHz for acquisition. The optics of the diagnostic is arranged such that the vertical antenna pattern forms a Gaussian beam of waist radius $\sim 3$ cm near the vessel mid-plane. In vacuum, the antenna pattern terminates, at the bottom of the machine, on a highly absorbing viewing dump [22], see Figure 1a. During plasma discharges, the antenna pattern can be shifted away from the viewing dump, due to refraction, at a threshold electron density close to $1\times 10^{19}\mathrm{m}^{-3}$[21]. The diagnostic measures fast electron emission at downshifted third or fourth harmonic emission as shown in Figure 1b. The electron energy that can be measured at each frequency is calculated in[19]. It is calculated that the maximum energy the diagnostic can measure without harmonic overlap is in the order of $\sim 250$ keV. The ECCD experiments discussed in this paper exploit the X2 gyrotron from the ECH, Electron Cyclotron Heating system on TCV[23]. The gyrotron produces wave at $82.7$ GHz and nominal power $\sim 680$ kW. The gyrotron is connected to a launcher which direct the beam into the plasma. Current drive is achieved by imparting a parallel component, with respect to the plasma magnetic field, to the ECH wave vector. The location of the beam power deposition along the radial coordinate is generally controlled by the chosen value of magnetic field strength. The toroidal and poloidal directions of the beam ((((in the reference frame of the plasma )))) are controlled by the launcher angles, so-called launcher toroidal angle, $\phi_{\mathrm{L}}$, and poloidal angle, $\theta_{\mathrm{L}}$ that can can be varied during discharges.
Figure 1: 1a V-ECE optics on TCV. The line of sight, which terminates on a beam dump (dark component below the machine floor) forms a Gaussian beam of radius $\sim$ 3 cm near the mid-plane. The radiation is reflected on an ellipsoidal mirror, split into X- and O- polarisations before detection in a heterodyne radiometer. 1b V-ECE frequency range, ECE harmonic frequencies ($f_{n\mathrm{ece}}$), and Right-Hand cut-off frequency ($f_{RH}$) on TCV with a toroidal magnetic field $B_{0}$=1.42 T at the vacuum vessel centre $R_{0}$=0.88 m. The represented plasma frequency is computed with a central electron density of 1.5 ×$\times 10^{19}$ m-3.
@@ 3.1 Background
The calibration of ECE diagnostics generally consists in determining, with the highest accuracy, the spectral response of the system to a given radiation power. Traditionally, two approaches have been retained for ECE calibration: (1) the use of sources of known characteristics as substitutes for the plasma and (2) the cross-calibration on a different, absolutely calibrated instrument. In the first approach, hot and cold sources are positioned within the vacuum chamber to irradiate the ECE antenna[24, 25]. The approach has been successfully employed since the 1970s in the vast majority of ECE systems in magnetic confinement devices, as demonstrated, for example, in Reference [26]. A limitation of this approach in the past was that the calibration needed to be conducted inside the vacuum vessel, necessarily during a machine vent. The system’s response, determined during a hot/cold source calibration, was typically used for several months, or even years, until the next calibration was feasible. This raised the question of the calibration’s stability in relation to variations in experimental conditions, which could occur on a daily basis. The limitations of the calibration with a reference source inside the vacuum vessel have been addressed in previous works, such as in Reference [27] by using external replica systems of ECE antennas, an application which is believed to be useful for ITER. The latter calibration approach exploits the absolute calibration of a different instrument within the same Tokamak, to obtain the response function of an ECE system. The ECE system can be cross-calibrated on another ECE system as in JET [28], or can be cross-calibrated on a different diagnostic which measures the same plasma parameters such as Thomson Scattering, inheriting all the uncertainties from the absolute calibration. This calibration method relies on the assumption that both diagnostics measure the same electron temperature. For ECE diagnostic systems, the measurement of the electron temperature requires that the plasma is optically thick and that the electron distribution is Maxwellian. When these conditions are fulfilled, successful cross-calibration of ECE systems with Thomson Scattering (TS) can be achieved, as demonstrated over the years, on TCV for example [29, 30, 31, 32, 33]. Historically, ECE and Thomson Scattering cross-calibration have used ohmic plasma discharges, which provide the best experimental plasma conditions allowing the assumption of a Maxwellian electron distribution. In other experimental conditions, such as those with NBI or ICRH heating, discrepancies between ECE and Thomson Scattering can occur, and modeling of the electron distribution is required [34, 35]. For a vertical ECE configuration, a direct cross-calibration with Thomson Scattering cannot be achieved for two main reasons. First, even in an optically thick plasma, the radiation intensity measured with a vertical line of sight cannot be straightforwardly associated with a local electron temperature, since the ECE resonance is not well localized in the vertical direction. Second, the plasma is not usually optically thick for a vertical ECE line of sight above a certain frequency (approximately 86 GHz on TCV). At these higher, optically thin frequencies, an accurate estimation of the absorption coefficient and, consequently, the radiation intensity needs to be performed in order to use the plasma emission for calibration. Examples of calculations of radiation intensity in optically thin conditions have been reported in previous works, such as in DIII-D for the measurement of the electron temperature from the optically gray third harmonic [36], and in JET for calibrating the optically thin O1 (first harmonic O-mode) radiation for temperature measurement in the inboard (High Field Side) pedestal [37]. In both cases, a horizontal lines of sight was employed.
@@ 3.4.1 Calculation of the peak power in a frequency band
The calibration factors are obtained by matching the modeled peak power with the measured voltage. The spectral intensities calculated for a few frequencies during discharge #$\#67946$ are shown in Figure 7 with their profile in each bandwidth. The parabolic shape of the intensity profile in the bandwidth justifies the need for optimization of the frequency step. The bandwidth power $P_{\mathrm{BW}}$, the total power in each frequency channel, is calculated as:
@@ 3.6 Discussing the uncertainties
The uncertainty in the V-ECE signal primarily arises from the digital (Butterworth) filtering of the raw data. Butterworth filtering is applied to the data to eliminate some noise in the detected signal, typically introduced by adjacent electronic components.
For the case of interest, the relative uncertainty on the filtered value used for the calibration is $\delta_{V}/{V}<5\%$. To assess the uncertainty in the modeled bandwidth power, the uncertainties in the parameters are propagated from the raw measurement to the calculated bandwidth power, as shown in in Figure 9. The assessment does not account for the potential uncertainty in the measurement of the magnetic field because the diagnostic is calibrated with thermal peak measurement, when the resonance is within the line of sight. The measured frequencies and the the line of sight constrain the value of the magnetic field at the resonance. The uncertainties that are propagated are those from the measurement of the electron density, $\delta_{n_{\mathrm{e}}}$ and electron temperature, $\delta_{T_{\mathrm{e}}}$. On TCV, The relative errors for both density and temperature measurements with Thomson Scattering are respectively below $5\%$ and $10\%$ in the plasma center. Both values increase significantly at the plasma edge, up to $25\%$ and $40\%$ respectively. The high uncertainties on the density and temperature profiles at the edge do not impact significantly the modeled intensity. That is because the optical depth of the plasma drops at the edge in such a way that the intensity depends mainly on the parameters in the plasma center.
@@ 4.1 ECCD measurement
Figure 11: Raw measurements of X-mode intensity from discharges #$\#73217$ and #$\#72644$ in the frequency range of $96-114$ GHz. The signals are plotted in decreasing frequency, from upper left to lower right. The energy measured at each frequency is calculated assuming a measurement of downshifted radiation from the third harmonic in both discharges. The subplot in the second row, second column, corresponds to the highest frequency and lowest energy (114 GHz, $\sim$76 keV in #$\#73217$ and $\sim$1 keV in #$\#72644$), while the subplot at the bottom right corresponds to the lowest frequency and highest energy (96 GHz, $\sim$176 keV in #$\#73217$ and $\sim$84 keV in #$\#72644$). The calculated energy differs between the two discharges due to the different values of magnetic field strength during the heating phase. The vertical dashed line represents the time when ECH heating starts in discharge #$\#72644$. The signal jump occurs farther from the vertical line at the highest energies.
@@
Calibrated V-ECE measurements during ECCD discharges #$\#73217$ and #$\#72644$ on TCV have been presented in Reference [19]. Raw measurements during those discharges are shown in Figure 11. Both discharges have similar values of ECH power. While the launcher toroidal angle is kept constant, the poloidal angle is varied in $5$ steps during the heating phase between $0.7$ s and $1.9$ s in #$\#72644$. In this configuration, as the EC beam propagates along the plasma equatorial mid-plane, the launcher angle controls the parallel component of the wave vector with respect to the toroidal magnetic field in the plasma. When this parallel wave-vector component is zero ($\phi_{\mathrm{L}}=0$), the EC wave purely heats the plasma by increasing the perpendicular energy of the resonant electrons around a given parallel velocity. Increasing the parallel wave vector (by increasing $\phi_{\mathrm{L}}$ ) results in a Doppler shift of the EC resonance, pulling a suprathermal electron tail out of the bulk at higher velocities than the corresponding heating case [42]. The variation of the poloidal angle of the launcher from $\sim 10^{\circ}$ to $\sim 26^{\circ}$ in each step allows the observation of a stair-shaped X-mode radiation intensity. When no heating power is applied in either plasma discharge, the measured radiation has similar intensity levels in both discharges, corresponding to the thermal background radiation originating from X2 emission. As soon as the heating is turned on in the discharge #$\#73217$ at $\sim 0.3$ s, all the channels immediately peak to reach a nearly stationary intensity level. We note how the intensity level during the discharge #$\#73217$ is generally higher than the intensity level in the discharge #$\#72644$ when the launcher configurations are similar. This is due to the more efficient current drive achieved in #$\#73217$, due in part to the lower density and higher electron temperature as the current drive efficiency is proportional, in general, to the ratio $T_{e}/n_{e}$ [43]. When the heating is turned on during the discharge #$\#72644$, at $\sim 0.7$ s, only the low energies (((($\sim$ $5-26$ keV)))) exhibit a sharp jump in the measured intensity. The intensity at the higher energies ($\sim$ $67-75$keV) shows a similar jump, with a delay of more than $\sim 500$ ms, when the launcher angle has been moved to its third stage. From that moment, the dynamics is inverted and the higher energies now have the sharpest jump in intensity compared to the lower energies. Interestingly, the measurement at the middle energies, in our range, will then reach the highest value of intensity when the launcher angle reaches its last stage. In measurements with Hard X-Ray tomographic Spectroscopy system (HXRS) [44] during the discharge #$\#72644$, the count rates shown in Figure 12 allow similar observations on the effects of the launcher angle sweep. In TCV, hard x-rays are emitted by suprathermal electron bremsstrahlung and are measured between 20 and 200 keV with an 8 keV resolution. In Figure 12, the sampling interval is set to 10 keV, which aligns with the resolution, making finer sampling unnecessary
@@
Simulations of the ECH wave interaction with the plasma, for the discharge #$\#72644$ is achieved using the the ray tracing code C3PO, solving the linear wave power absorption [45]. The refractive indices and the angle between the wave vector and the magnetic field, calculated at the wave absorption location for the various launcher angles are shown in Table 1. The increase in the launcher poloidal angle allows to achieve higher values ((((in absolute terms)))) of the parallel refractive index. Higher parallel refractive indices generally allow the intersection between the EC resonant curves ($v_{\perp}$ vs $v_{\parallel}$) and the $v_{\parallel}$ axis at higher values of $v_{\parallel}$ [46], see Figure 13. The physical mechanisms of the plasma-wave interaction mechanisms include, in general, momentum diffusion and pitch-angle scattering, which make the ECCD resonant interaction region complex to determine. For the purposes of this experiment it is observed that higher values $\theta_{L}$ and thus higher absolute values of $N_{\parallel}$ induce more energetic electrons in the plasma, explaining the different jumps in the measured intensities.
@@
Figure 14 show the jumps in intensity, relative to the thermal phase, measured at each frequency in the different stages of discharge #$\#72644$. The jump is obtained by dividing the intensity in each heated stage by the intensity in the thermal phase. It is seen how at relatively low parallel refractive index ($\left|N_{\parallel}\right|=0.22$ ), the highest frequencies ((((lowest energies)))) jump is the most important. As $\left|N_{\parallel}\right|$ increases, an intermediate frequency (intermediate energy as well), near $104$ GHz shows the sharpest jumps. We note that V-ECE X-mode intensity in ideal conditions depends on the density of the non-thermal electrons in a given energy band. That is, the highest jump observed around $104$ GHz, with increasing $\left|N_{\parallel}\right|$ suggests an increase in the number density of non-thermal electrons at the corresponding energy. This does not necessarily mean that the ECH wave is damped on electrons whose V-ECE energies are measured at $104$ GHz when the refractive index increases. The increase of the number density at the given energy, which explains the sharp jump, is a consequence of the overall dynamics in the electron energy space, originally caused by the ECH wave.
@@
First-principle calculation using the code LUKE is used compute the electron energy distribution at the different stages of discharge #$\#72644$. The first calculation is shown in Figure 15a. The total number of electrons above a certain energy is plotted for a Maxwellian distribution and for the computed distribution. At a given energy, say 15 keV, which is above the electron thermal energy, the total number of electrons above that energy increases over time intervals (with increasing $\left|N_{\parallel}\right|$). This results confirms that low-energy electrons are pushed into the fast-electrons tail with ECCD and the number of transported electrons increases with parallel refractive index. Figure 15b shows the derivative of the number of electrons in the modeled distribution deduced from the density in the Maxwellian distribution. The discontinuity in the spectra is due to the logarithmic scale and to the absolute value in the plotted quantity. It represents the regions where the density in the Maxwellian distribution changes from higher to lower than that of the modeled distribution, which is the region in the phase space where the electrons are being pumped out by the ECH wave.